OpenStudy (anonymous):

6th root of 7 times the square root of 21

OpenStudy (anonymous):

\[\sqrt[6]{7}\times \sqrt[2]{21}\]

OpenStudy (anonymous):

6.33810341019

OpenStudy (anonymous):

here are the answers \[\sqrt[8]{21}\] \[\sqrt[12]{3}\] \[\sqrt[12]{21}\] \[\sqrt[84]{3}\]

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@cheylove32 i figured that out, but the problem doesnt have that as an answer, sorry

OpenStudy (anonymous):

thanks anyways

OpenStudy (anonymous):

Okay sorry!

OpenStudy (solomonzelman):

\[\Large \sqrt[6]{7} \times \sqrt[2]{21}\]\[\Large\sqrt[6]{7} \times \sqrt[2]{21^1}\]\[\Large \sqrt[6]{7} \times \sqrt[2\times 3]{21^{1\times 3}}\]\[\Large \sqrt[6]{7} \times \sqrt[6]{21^{3}}\]\[\Large \sqrt[6]{7 \times 21^{3}} \]\[\Large \sqrt[6]{7 \times9261} \]\[\Large \sqrt[6]{64827} \]

OpenStudy (solomonzelman):

This is the final result.

OpenStudy (anonymous):

thanks anyways but i figured it out

OpenStudy (anonymous):

it was D